# Long-term stable nonlinear evolutions of ultracompact black-hole mimickers

### v1

@@ I Introduction
Black-hole solutions in closed analytic form have played a crucial role in many aspects of the theoretical understanding of general relativity (GR) almost right from the moment Einstein published his theory. Starting with observations of X-ray sources (e.g. Cygnus X-1 Bowyer et al. (1965)) and quasars (e.g. 3C 273 Hazard et al. (1963); Schmidt (1963)), and their eventual identification with accretion onto compact objects, black holes (BHs) acquired an equally prominent role as seemingly common citizens of our universe. The rapidly increasing number of gravitational-wave (GW) detections by LIGO and Virgo Abbott et al. (2016, 2023), all compatible with theoretical predictions for BHs and neutron stars in GR Abbott et al. (2021), and direct observations of BH shadows by the Event Horizon Telescope Akiyama et al. (2019, 2022) has further cemented the prominence of BHs in observational astronomy.
@@
In both mathematical modeling as well as popular imagination, the most characteristic feature of a BH is its event horizon, the boundary of an interior region from which nothing, not even light, can escape. In terms of their observational phenomenology, however, the decisive feature of BHs are their (unstable) light rings (LRs), circular null-geodesics on which photons can orbit; for a Schwarzschild BH of mass $M$, the event horizon is located at areal radius $r=2GM/c^{2}$ and the LR at $r=3GM/c^{2}$. The shape of BH shadows, accretion dynamics and the BH ringdown spectrum are closely related to LRs and the perturbational potential in their vicinity Synge (1966); Cardoso et al. (2009); Koga et al. (2022); Völkel et al. (2022). The observational evidence therefore points more directly at the existence of LRs, not necessarily BHs. Additionally, classical black holes come with theoretical problems, like the singularity and the information paradox Mathur (2009), motivating the study of BH mimickers. The GW emission from compact boson-star (BS) binaries furthermore exhibits significant degeneracy with that from BH binaries Evstafyeva et al. (2024). Even if one accepts Occam’s razor and parsimoniously interprets BHs as the most likely sources, there remains the intriguing possibility that a population of ultra-compact objects (UCOs) is hiding inside a more conventional BH society and may provide us with clues about dark-matter candidate fields they are composed of.
@@
Clearly, dynamical longevity is a key requirement of any viable alternative to the BH paradigm. The stability of UCOs or BH mimickers – both implied henceforth to denote objects with a LR but no horizon – has, consequently, gained much attention in recent years. Under a broad set of assumptions, BH mimickers have been shown to possess two LRs, one stable and one unstable located at the minimum and maximum, respectively, of the null geodesic’s potential Cunha et al. (2017); Cunha and Herdeiro (2020); Kerr BHs, in contrast only have unstable LRs. The existence of a minimum in the effective potential leads to stable trapping, i.e. null geodesics originating close to the minimum may remain indefinitely bound to it. At linear level, waves propagating on a background spacetime with stable trapping exhibit decay slower than polynomial or exponential as occurs for unstable trapping Keir (2016). Guided by results on the instability of anti-de Sitter spacetime and its similar decay rates Bizoń and Rostworowski (2011); Holzegel and Smulevici (2013, 2014), it has been heuristically argued that UCOs may also be nonlinearly unstable, thereby ruling them out as plausible BH mimickers Cardoso et al. (2014). In the absence of rigorous analytic results, however, numerical evolutions are the only available tool for testing this hypothesis.
@@
To date, literature results investigating the ultimate fate and stability of UCOs are inconclusive. By simulating two families of ultracompact BSs, Ref. Cunha et al. (2023) either finds migration to nonultracompact configurations or collapse to a BH, typically on timescales $\lesssim\mathcal{O}(10^{3})\,\mu^{-1}$ in units of the scalar mass parameter, and concludes that the LRs trigger instability. Reference Cunha (2025) explains the instability as the result of the backreaction of perturbations at the stable LR appearing to deepen the geodesic potential. In contrast, simulations of the merger and ringdown of ultracompact spinning BSs yield no evidence of this instability Siemonsen (2024), although the timescales the simulations span in that work are considerably shorter. Further simulations in spherical symmetry of scalarized Reisner-Nordström BHs with LRs also exhibited long-term stability Guo et al. (2024). In a recent study using a toy-model spacetime with two LRs, Ref. Benomio et al. (2024) analyzes the differences between linear and non-linear dynamics in more detail: within the region of stable trapping, the slow linear decay leads to a growth of higher-order derivatives and energy in the nonlinear evolution, thereby resulting in turbulent behaviour. This phenomenon is proven in the perturbative regime Redondo-Yuste and Cárdenas-Avendaño (2025), where the authors argue how it may not necessarily trigger an instability. The ultimate end state of these configurations, however, remains an open question.
@@
Here, we explore the stability of a class of spherically symmetric ultracompact BSs with LRs, characterized by thin-shell profiles Collodel and Doneva (2022) and reaching compactness $C\sim 0.32$. We combine a set of numerical simulations in spherical symmetry (1+1), axisymmetry (2+1) and full 3+1 dimensions with perturbative calculations of the radial oscillation frequency, observing excellent agreement between all approaches compatible with the standard result that stable and unstable branches of stars, including UCOs with LRs, are separated by local extrema in the mass-radius diagram. In particular, we observe no signs of instability on the stable branch containing BH mimickers in agreement with the restabilization, i.e. real frequency, of the fundamental radial oscillation mode predicted in this regime by our perturbative calculations.
@@
The paper is organised as follows. Sec. II presents the BS models considered, as well as the different codes. Sec. III summarizes our evolutions, showcasing their stability. We connect the observed dynamics to linear perturbation theory in Sec. IV, and demonstrate that the light ring structure is preserved in Sec. V. We supplement our detailed evolutions with a wider range of models in Sec. VI, before concluding in Sec. VII. Henceforth, we employ Planck units $c=G=\hbar=1$, but keep the scalar-field mass parameter $\mu$ which controls the scale invariance of the spacetimes in question. Code units are translated into SI (or similar) by fixing $\mu$ and scaling with appropriate powers thereof: e.g. $\mu=1.33\times 10^{-10}\,{\rm eV}$ corresponds to $\mu^{-1}=M_{\odot}=1.48\,{\rm km}$.
II Theory and numerical codes
@@
We consider BSs in GR composed of one complex scalar field $\varphi$ with solitonic potential $V(\varphi)$ minimally coupled to gravity by the action
@@
We compute stationary BS models in spherical symmetry writing $\varphi(t,r)=A(r)e^{\mathrm{i}\mkern 1.0mu\omega t}$ and using a GR version with quadruple-precision of the two-way shooting code of Ref. Evstafyeva et al. (2023). The resulting single-BS spacetimes are evolved in time using four different codes in spherical symmetry, axisymmetry or full $3D$ using the Baumgarte-Shapiro-Shibata-Nakamura (BSSN) Baumgarte and Shapiro (1998); Shibata and Nakamura (1995) and the conformal covariant $Z4$ (CCZ4) Alic et al. (2012) formulations. (i) sphericalbsevolver achieves spherically symmetry time evolutions through dimensional reduction with $SO(3)$ symmetry according to Ref. Cook et al. (2016) using a uniform radial grid. (ii) A $2D$ code implements axisymmetry using the same modified cartoon method but for $SO(2)$ isometry. It is based on (iii) ExoZvezda exo (2024), a code for exotic matter built on GRChombo Radia et al. (2022); Andrade et al. (2021), which evolves $3D$ domains with adaptive mesh refinement provided by chombo Adams et al. (2019). (iv) lean Sperhake (2007) is based on the cactus computational toolkit Allen, G. and Goodale, T. and Massó, J. and Seidel, E. (1999) and evolves $3D$ domains with mesh refinement provided by Carpet Schnetter et al. (2004). In the following, we will refer to these codes and their respective simulations with the acronyms Sph, Axi, E3D and L3D. Unless specified otherwise, our simulations employ a resolution $h=1/(12\mu)$ with outer boundary at $256\,\mu^{-1}$ on a unigrid (Sph) or using 5 refinement levels (Axi, E3D, L3D).
III Time evolutions
Our study focuses on BS families for two values of the solitonic parameter, $\sigma_{0}=0.08$ and $\sigma_{0}=0.06$; the resulting BS models are displayed in the plane spanned by the central scalar amplitude $A_{0}$ and the ADM mass $M$ in Fig. 1. There we mark in orange color the ultracompact BSs posessing a pair of LRs; for both $\sigma_{0}$ values, this set of UCOs extends to the left, beyond the global maximum of the $M(A_{0})$ curve, i.e. into the radially stable branch. This regime comprises the stars of interest in our study and we mark for each potential one representative specimen with properties as summarized in Table 1.
Figure 1: The ADM mass is shown versus the central scalar-field amplitude $A_{0}$ for the one-parameter families of BS models with solitonic potential parameter $\sigma_{0}=0.06$ (black) and $\sigma_{0}=0.08$ (blue). For each family solid (dashed) segments mark perturbatively stable (unstable) branches and we highlight in orange color BSs exhibiting a pair of LRs. The insets indicate the perturbatively stable BH mimickers S06A044 and S08A06 that form the focus of our numerical investigation.
Table 1: Nickname, solitonic potential parameter $\sigma_{0}$, central scalar-field amplitude $A_{0}$, ADM mass $M$, Noether charge $\mathcal{N}$, the radius $r_{99}$ encompassing $99\,\%$ of the total mass, compactness $C$ and fundamental radial oscillation frequency $\chi_{0}^{2}$ for the two BS models marked by square and circle in Fig. 1.
@@ III Time evolutions
Figure 2: Time evolution of the deviation $\Delta A_{0}$ of the central scalar-field amplitude for the S06A044 (top) and S08A06 models (bottom) obtained in spherical symmetry, axi-symmetry and $3D$. The linear drift visible for S08A06 converges away as shown in the inset. The four symbols in the top panel mark times when we evaluate the LR potential; cf. Fig. 4.
@@
We begin our assessment of these two models by plotting in Fig. 2 the deviation $\Delta A_{0}$ of the central scalar-field amplitude from its initial value over time as obtained with different symmetry assumptions. For the S06A044 model, we observe periodic oscillations with a relative magnitude $\sim\!\!0.3\,\%$ and good agreement by all codes. For S08A06, the oscillations are two orders of magnitude smaller and exhbit a linear drift which converges away with increasing resolution. This is demonstrated by the three blue-shaded curves obtained with L3D using different grid spacing $h$. Closer analysis of the differences yields $5^{\rm th}$-order convergence and the Richardson extrapolated function oscillates with high precision around zero.
@@
Two features manifest in this figure require further explanation. First, we attribute the considerably larger oscillations observed for S06A044 to a larger numerical error in the initial data; for this model, round-off uncertainty in the BS frequency $\sim\!\!10^{-32}$ can effect relative variations in the ADM mass or central lapse comparable to the level $\Delta A_{0}/A_{0}$ observed. The S08A06 model is less challenging and results in initial data with small uncertainty compared to those incurred in the time evolutions. This feature also explains the level of agreement between time evolutions employing different symmetry: the perturbations of S06A044 are initial-data dominated with comparatively smaller contributions from the time evolution. For S08A06, on the other hand, details of the time evolution dominate, resulting in more pronounced relative – albeit tiny absolute – differences. This observation is confirmed by artificially lowering the quality of the initial data or time evolution which amplify drift or oscillation magnitude but with no change in frequency. The $3D$ simulations displayed here have been computed using octant symmetry. We have performed additional versions of the $3D$ simulations of S08A06 using L3D (E3D) with $h=1/12$ in full grid mode up to $t\gtrsim 3000$ ($t\gtrsim 1000$); these give identical results to the octant runs within round-off precision.
@@
The time evolution of the maximal Kretschmann scalar $R^{\alpha\beta\gamma\delta}R_{\alpha\beta\gamma\delta}$ exhibits, up to an overall scaling, nearly identical behaviour as $\Delta A_{0}$, including the oscillatory pattern, drift and $5^{\rm th}$-order convergence towards the equilibrium value. The Noether charge is conserved with identical convergence to within $\Delta\mathcal{N}/\mathcal{N}\lesssim 10^{-3}$ and $10^{-4}$ for S06A044 and S08A06, respectively, predicting $\Delta\mathcal{N}/\mathcal{N}\lesssim 10^{-6}$ in the continuum limit for S08A06. The constraints exhibit a rapid drop to $L_{2}$ norms $\sim 10^{-9}$ due to the CCZ4 damping Alic et al. (2013), but converge only between $1^{\rm st}$ and $4^{\rm th}$ order.
IV Perturbative predictions
The oscillatory behaviour of $\Delta A_{0}$ in Fig. 2 can be understood in terms of linearized calculations. For this purpose, we compute the frequency $\chi_{0}$ of the fundamental radial oscillation mode for our BS families using the approach of Refs. Gleiser and Watkins (1989); Kain (2021); see App. C for details. The transitions between stable and unstable branches in Fig. 1 are characterized by zero crossings of $\chi_{0}^{2}$ such that a real (imaginary) $\chi_{0}$ implies radial (in)stability. For our BS models S06A044 and S08A06, we obtain $\chi_{0}^{2}>0$ as expected from their location on the rising slope of the $M(A_{0})$ curve. The behaviour of these BSs, as shown in Fig. 2, exhibits excellent agreement with the perturbative predictions: numerical imperfections trigger perturbations around the equilibrium solution whose time evolution is dominated by radial oscillations with the expected frequency. We quantify this agreement by plotting in Fig. 3 for S08A06 the Fourier spectrum of $\Delta A_{0}(t)$, i.e. the data of Fig. 2. The spectrum peaks at the $\chi_{0}$ value predicted by the perturbative calculation. The corresponding analysis for S06A044 yields similar agreement (see App. B).
@@ IV Perturbative predictions
V Light-ring structure
Having demonstrated the long-term stable behaviour of our BH mimickers S06A044 and S08A06, we now turn to the question whether their LR structure remains preserved. In a stationary and axisymmetric spacetime, a LR is defined as a null geodesic with a tangent vector field that is always a linear combination of the Killing vectors $\partial_{t}$ and $\partial_{\phi}$, assuming coordinates $(t,r,\theta,\phi)$ adapted to the symmetry. The conditions for a LR can then be recast as Cunha et al. (2017)
$\partial_{\mu}H_{\pm}=0\,,~{}~{}~{}H_{\pm}=\frac{-g_{t\phi}\pm\sqrt{g_{t\phi}^{2}-g_{tt}g_{\phi\phi}}}{g_{\phi\phi}}\,,$ (3)
and (un)stable LRs are located at local (maxima) minima of $H_{\pm}$. In spherical symmetry with adapted coordinates $H_{\pm}$ simplifies to a single potential function $H=\sqrt{-g_{tt}/g_{\phi\phi}}$.
[…]
@@ V Light-ring structure
Figure 4: (left) Areal radius of the (un)stable light ring over time, as extracted from the AEP $H_{\rm eff}$ (right) during the time evolution of S06A044 performed with E3D. The inset shows the difference between the extrema of the AEP over time. The green markers correspond to the snapshots of the AEP given in the right panel. The grey lines indicate the positions of the extrema. The solid line is the effective potential (3) at the start of the simulation.
@@
Our analysis of the resulting potentials for S06A044 and S08A06 indeed yields one minimum and one maximum at all $t$. The location of the respective LRs exhibits minor periodic variation commensurate with the BS oscillations observed in the other diagnostics, but the pair of light rings LRs remains in place with no indication of instability. We show these variations together with selected profiles of the AEP over a representative oscillation cycle in Fig. 4 for S06A044. The same behaviour, merely with smaller oscillation amplitude, is found for S08A06.
VI Non-linear stability
We complement our detailed analysis of the BS models S06A044 and S08A06 with a larger set of time evolutions of the $\sigma=0.08$ BS sequence. We have used all codes for this analysis but restrict $3D$ evolutions to $\mu t\lesssim 3000$ for reasons of computational cost. In this analysis we define the instability time as the moment when $A_{0}(t)$ departs by $10\,\%$ from its equilibrium value. If $A_{0}(t)$ yields no signature of departing from the oscillatory pattern seen in Fig. 2 within $3000\,\mu^{-1}$, the model is designated stable. The result is shown in Fig. 5 and fully agrees with the perturbative predictions: models on unstable branches succumb to exponentially growing perturbations on a time scale that diverges at the threshold to stable branches. These transition points coincide with extrema in the $M(A_{0})$ curves, i.e. points where the oscillation frequency $\chi_{0}$ of the fundamental mode switches between real and imaginary.
@@ VI Non-linear stability
Figure 5: Instability timescales as a function of $A_{0}$ for solitonic BSs with $\sigma_{0}=0.08$. The vertical lines mark the transition between stable and unstable branches as predicted by linear theory; cf. Fig. 1. Marker style indicates dynamical fate while colour indicates code used: black is Sph, blue Axi, orange L3D. BS models whose time evolution show no sign of instability are represented with arrows at the top of the figure.
VII Concluding remarks

### v2

@@ Abstract
Introduction—Black-hole solutions in closed analytic form have played a crucial role in many aspects of the theoretical understanding of general relativity (GR) almost right from the moment Einstein published his theory. Starting with observations of X-ray sources (e.g. Cygnus X-1 Bowyer et al. (1965)) and quasars (e.g. 3C 273 Hazard et al. (1963); Schmidt (1963)) and their eventual identification with accretion onto compact objects, black holes (BHs) acquired an equally prominent role as seemingly common citizens of our Universe. The rapidly increasing number of gravitational-wave (GW) detections by LIGO and Virgo Abbott et al. (2016, 2023), all compatible with theoretical predictions for BHs and neutron stars in GR Abbott et al. (2021), and direct observations of BH environments by the Event Horizon Telescope Akiyama et al. (2019, 2022) have further cemented the prominence of BHs in observational astronomy.
@@
In both mathematical modeling as well as popular imagination, the most characteristic feature of a BH is its event horizon, the 2+1 dimensional outer boundary of the set of spacetime points from which no null geodesics can reach future null infinity. In terms of their observational phenomenology, however, the decisive feature of BHs is their (unstable) light rings (LRs), circular null geodesics on which photons can orbit; for a Schwarzschild BH of mass $M$, the event horizon is located at areal radius $r=2GM/c^{2}$ and the LR at $r=3GM/c^{2}$. The appearance of accreting BHs is strongly influenced by the presence and properties of the LR, while quasinormal modes have been shown to have a close connection with null unstable geodesics Synge (1966); Cardoso et al. (2009); Jusufi (2020); Koga et al. (2022); Völkel et al. (2022); Pedrotti and Calzà (2025). The observational evidence therefore points more directly at the existence of LRs, not necessarily BHs. Additionally, classical black holes come with theoretical issues, like the singularity problem Hawking and Ellis (1973) or the information paradox Mathur (2009), motivating the study of BH mimickers. Even if one accepts Occam’s razor and parsimoniously interprets BHs as the most likely sources, there remains the intriguing possibility that a population of ultracompact objects (UCOs) is hiding inside a more conventional BH society and may provide us with clues about the dark-matter candidate fields they are composed of. We also note that the GW emission from compact boson-star (BS) binaries exhibits significant degeneracy with that from BH binaries Evstafyeva et al. (2024).
@@
Clearly, dynamical longevity is a key requirement of any viable alternative to the BH paradigm. The stability of UCOs or BH mimickers – both implied henceforth to denote objects with a LR but no horizon – has, consequently, gained much attention in recent years. Under a broad set of assumptions, BH mimickers have been shown to possess two LRs, one stable and one unstable located at the minimum and maximum, respectively, of the null geodesic’s potential Cunha et al. (2017); Cunha and Herdeiro (2020); Kerr BHs, in contrast have unstable LRs only. The existence of a minimum in the effective potential leads to stable trapping, i.e. null geodesics originating close to the minimum may remain indefinitely bound to it. At linear level, waves propagating on a background spacetime with stable trapping exhibit decay slower than polynomial or exponential as occurs for unstable trapping Keir (2016). Guided by results on the instability of anti-de Sitter spacetime and its similar decay rates Bizoń and Rostworowski (2011); Holzegel and Smulevici (2013, 2014), it has been heuristically argued that UCOs may also be nonlinearly unstable, thus ruling them out as plausible BH mimickers Cardoso et al. (2014). In the absence of rigorous analytic results, however, numerical evolutions are the only available tool for testing this hypothesis.
@@
To date, literature results investigating the ultimate fate and stability of UCOs are inconclusive. By simulating two families of ultracompact BSs, Ref. Cunha et al. (2023) finds either migration to nonultracompact configurations or collapse to a BH, typically on timescales $\lesssim\mathcal{O}(10^{3})\,\mu^{-1}$ in units of the scalar mass parameter, and concludes that the LRs trigger instability. Reference Cunha (2025) explains the instability as the result of the backreaction of perturbations at the stable LR appearing to deepen the geodesic potential. In contrast, simulations of the merger and ringdown of ultracompact spinning BSs yield no evidence of this instability Siemonsen (2024), although the timescales the simulations span in that work are considerably shorter. Further simulations in spherical symmetry of scalarized Reisner-Nordström BHs with LRs also exhibited long-term stability Guo et al. (2024). In a recent study using a toy-model spacetime with two LRs, Ref. Benomio et al. (2024) analyzes the differences between linear and noninear dynamics in more detail: within the region of stable trapping, the slow linear decay leads to a growth of higher-order derivatives and energy in the nonlinear evolution, thereby resulting in turbulent behaviour. This phenomenon is proven in the perturbative regime Redondo-Yuste and Cárdenas-Avendaño (2025), where the authors argue how it may not necessarily trigger an instability. The ultimate end state of these configurations, however, remains an open question.
@@
Here, we explore the stability of a class of spherically symmetric ultracompact BSs with LRs characterized by thin-shell profiles Collodel and Doneva (2022) and reaching compactness $C\sim 0.32$. We combine a set of numerical simulations in spherical symmetry (1+1), axisymmetry (2+1), and full 3+1 dimensions with perturbative calculations of the radial oscillation frequency, observing excellent agreement between all approaches compatible with the standard result that stable and unstable branches of stars, including UCOs with LRs, are separated by local extrema in the mass-radius diagram. In particular, we observe no signs of instability on the stable branch containing BH mimickers, in agreement with the restabilization, i.e. real frequency, of the fundamental radial oscillation mode predicted in this regime by our perturbative calculations.
Figure 1: The ADM mass is shown versus the central scalar-field amplitude $A_{0}$ for the one-parameter families of BS models with solitonic potential parameter $\sigma_{0}=0.06$ (black) and $\sigma_{0}=0.08$ (blue). For each family, solid (dashed) segments mark perturbatively stable (unstable) branches and we highlight in orange color BSs exhibiting a pair of LRs. The insets indicate the two perturbatively stable BH mimickers with parameters detailed in Table 1 that form the focus of our numerical investigation.
@@
Henceforth, we employ Planck units $c=G=\hbar=1$ but keep the scalar-field mass parameter $\mu$ which controls the scale invariance of the spacetimes in question. Code units are translated into SI (or similar) by fixing $\mu$ and scaling with appropriate powers thereof: e.g. $\mu=1.33\times 10^{-10}\,{\rm eV}$ corresponds to $\mu^{-1}=M_{\odot}=1.48\,{\rm km}$.
@@
Theory and numerical codes—We consider BSs in GR composed of one minimally coupled, complex scalar field $\varphi$, with solitonic potential, described by the action
@@
We compute stationary BS models in spherical symmetry writing $\varphi(t,r)=A(r)e^{\mathrm{i}\mkern 1.0mu\omega t}$ and using a GR version with quadruple precision of the two-way shooting code of Ref. Evstafyeva et al. (2023). The resulting single-BS spacetimes are evolved in time using four different codes in spherical symmetry, axisymmetry, or full 3D using the Baumgarte-Shapiro-Shibata-Nakamura (BSSN) Baumgarte and Shapiro (1998); Shibata and Nakamura (1995) and the conformal covariant $Z4$ (CCZ4) Alic et al. (2012) formulations. (i) sphericalbsevolver achieves spherically symmetric time evolutions through dimensional reduction with $SO(3)$ symmetry according to Ref. Cook et al. (2016) using a uniform radial grid. (ii) A 2D code implements axisymmetry using the same modified cartoon method but for $SO(2)$ isometry. It is based on (iii) ExoZvezda new (2024), a code for exotic matter built on GRChombo Radia et al. (2022); Andrade et al. (2021), which evolves 3D domains with adaptive mesh refinement provided by chombo Adams et al. (2019). (iv) lean Sperhake (2007) is based on the cactus computational toolkit Allen, G. and Goodale, T. and Massó, J. and Seidel, E. (1999) and evolves 3D domains with mesh refinement provided by Carpet Schnetter et al. (2004). We will refer to these codes and their respective simulations with the acronyms Sph, Axi, E3D and L3D. Unless specified otherwise, our simulations employ $4^{\rm th}$-order discretization, a resolution $h=1/(12\mu)$ with outer boundary at $256\,\mu^{-1}$ on a unigrid (Sph) or using 5 refinement levels (Axi, E3D, L3D).
Table 1: Nickname, solitonic parameter $\sigma_{0}$, central scalar-field amplitude $A_{0}$, ADM mass $M$, Noether charge $\mathcal{N}$, the radius $r_{99}$ encompassing $99\,\%$ of the total mass, compactness $C\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=M/r_{99}$ and fundamental radial oscillation frequency $\chi_{0}^{2}$ for the two BS models marked by square and circle in Fig. 1.
@@
Figure 2: Time evolution of the deviation $\Delta A_{0}$ of the central scalar-field amplitude for the S06A044 (top) and S08A06 (bottom) models obtained in spherical symmetry, axisymmetry and 3D. The linear drift visible for S08A06 converges away as shown in the inset. The four symbols in the top panel mark times when we evaluate the LR potential; cf. Fig. 4.
Time evolutionsOur study focuses on BS families for two values of the solitonic parameter, $\sigma_{0}=0.08$ and $\sigma_{0}=0.06$; the resulting BS models are displayed in the plane spanned by the central scalar amplitude $A_{0}$ and the Arnowitt-Deser-Misner (ADM) mass $M$ in Fig. 1. There, we mark in orange color the ultracompact BSs possessing a pair of LRs; for both $\sigma_{0}$ values, this set of UCOs extends to the left, beyond the global maximum of the $M(A_{0})$ curve, i.e. into the radially stable branch. This regime comprises the stars of interest in our study and we mark for each potential one representative specimen with properties as summarized in Table 1.
@@
We begin our assessment of these two models by plotting in Fig. 2 the deviation $\Delta A_{0}$ of the central scalar-field amplitude from its initial value over time as obtained with different symmetry assumptions. For the S06A044 model, we observe periodic oscillations with a relative magnitude $\sim\!\!\!0.3\,\%$ and good agreement by all codes. For S08A06, the oscillations are 2 orders of magnitude smaller and exhibit a linear drift which converges away with increasing resolution. This is demonstrated by the three blue-shaded curves obtained with L3D using different grid spacing $h$. Closer analysis of the differences yields $5^{\rm th}$-order convergence and the Richardson extrapolated function oscillates with high precision around zero.
@@
Two features manifest in this figure require further explanation. First, we attribute the considerably larger oscillations observed for S06A044 to a larger numerical error in the initial data; for this model, round-off uncertainty in the BS frequency $\sim\!\!10^{-32}$ can effect relative variations in the ADM mass or central lapse comparable to the level $\Delta A_{0}/A_{0}$ observed. The S08A06 model is less challenging and results in initial data with small uncertainty compared to those incurred in the time evolutions. This feature also explains the level of agreement between time evolutions employing different symmetry: the perturbations of S06A044 are initial-data dominated with comparatively smaller contributions from the time evolution. For S08A06, on the other hand, details of the time evolution dominate, resulting in more pronounced relative – albeit tiny absolute – differences. This observation is confirmed by artificially lowering the quality of the initial data or time evolution which amplifies drift or oscillation magnitude but with no change in frequency. The 3D simulations displayed here have been computed using octant symmetry. We have performed additional versions of the 3D simulations of S08A06 using L3D (E3D) with $h=1/12$ in full grid mode up to $t\gtrsim 3000$ ($t\gtrsim 1000$); these give identical results to the octant runs within round-off precision.
@@
The time evolution of the maximal Kretschmann scalar $R^{\alpha\beta\gamma\delta}R_{\alpha\beta\gamma\delta}$ exhibits, up to an overall scaling, nearly identical behaviour as $\Delta A_{0}$, including the oscillatory pattern, drift and $5^{\rm th}$-order convergence toward the equilibrium value. The Noether charge is conserved with identical convergence to within $\Delta\mathcal{N}/\mathcal{N}\lesssim 10^{-3}$ and $10^{-4}$ for S06A044 and S08A06, respectively, predicting $\Delta\mathcal{N}/\mathcal{N}\lesssim 10^{-6}$ in the continuum limit for S08A06. The constraints exhibit a rapid drop to $L_{2}$ norms $\sim 10^{-9}$ due to the CCZ4 damping Alic et al. (2013), but converge only between $1^{\rm st}$ and $4^{\rm th}$ order; cf. Fig. S3 in the Supplemental Material.
@@
Perturbative predictions—The oscillatory behaviour of $\Delta A_{0}$ in Fig. 2 can be understood in terms of linearized calculations. For this purpose, we compute the frequency $\chi_{0}$ of the fundamental radial oscillation mode for our BS families using the approach of Refs. Gleiser and Watkins (1989); Kain (2021); cf. Sec. III in the Supplemental Material. The transitions between stable and unstable branches in Fig. 1 are characterized by zero crossings of $\chi_{0}^{2}$ such that a real (imaginary) $\chi_{0}$ implies radial (in)stability. For our BS models S06A044 and S08A06, we obtain $\chi_{0}^{2}>0$ as expected from their location on the rising slope of the $M(A_{0})$ curve. The behaviour of these BSs, as shown in Fig. 2, exhibits excellent agreement with the perturbative predictions: numerical imperfections trigger perturbations around the equilibrium solution whose time evolution is dominated by radial oscillations with the expected frequency. We quantify this agreement by plotting in Fig. 3 for S08A06 the Fourier spectrum of $\Delta A_{0}(t)$, i.e. the data of Fig. 2. The spectrum peaks at the $\chi_{0}$ value predicted by the perturbative calculation. The corresponding analysis for S06A044 yields similar agreement; cf. Fig. S4 in the Supplemental Material.
@@
Figure 4: (left) Areal radius of the (un)stable light ring over time as extracted from the AEP $H_{\rm eff}$ (right) during the time evolution of S06A044 performed with E3D. The inset shows the difference between the extrema of the AEP over time. The green markers correspond to the snapshots of the AEP given in the right panel. The gray lines indicate the positions of the extrema. The solid line is the effective potential (3) at the start of the simulation.
Light-ring structure—Having demonstrated the long-term stable behaviour of our BH mimickers S06A044 and S08A06, we now turn to the question whether their LR structure remains preserved. In a stationary and axisymmetric spacetime, a LR is defined as a null geodesic with a tangent vector field that is always a linear combination of the Killing vectors $\partial_{t}$ and $\partial_{\phi}$, assuming coordinates $(t,r,\theta,\phi)$ adapted to the symmetry. The conditions for a LR can then be recast as follows [cf. Eqs. (1)-(8) in Ref. Cunha et al. (2017)]:
$\partial_{\mu}H_{\pm}=0\,,~~\text{where}~~H_{\pm}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\frac{-g_{t\phi}\pm\sqrt{g_{t\phi}^{2}-g_{tt}g_{\phi\phi}}}{g_{\phi\phi}}\,.$ (3)
Furthermore, (un)stable LRs are located at local (maxima) minima of $H_{\pm}$. In spherical symmetry with adapted coordinates, $H_{\pm}$ simplifies to a single potential function $H=\sqrt{-g_{tt}/g_{\phi\phi}}$.
Under the assumption that an evolved spacetime remains nearly spherically symmetric and stationary, the potential function can be approximated by an adiabatic effective potential (AEP) $H_{\rm eff}$ computed via an averaging procedure analogous to Ref. Cunha et al. (2023). Specifically, we transform between standard Cartesian and spherical coordinates $x^{\alpha}=(t,x,y,z)$, $\hat{x}^{\mu}=(t,R,\theta,\phi)$. For a sphere of fixed coordinate radius $R$, we compute the proper area $\mathcal{A}$, the areal radius $r\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\sqrt{\mathcal{A}/(4\pi)}$, and the AEP $H_{\rm eff}(r)=\sqrt{(\langle\alpha^{2}\rangle-\langle\gamma_{ij}\beta^{i}\beta^{j}\rangle)}/r$, where $\alpha$, $\beta^{i}$ and $\gamma_{ij}$ are the lapse, shift and spatial metric in Cartesian coordinates, and $\langle\,.\,\rangle$ denotes averaging over the sphere. Repeating this procedure for a sequence of $\sim 50$ coordinate spheres with $\Delta R\approx 0.1$ across the relevant shell region, we obtain at each time $t$ an estimate for the potential $H_{\rm eff}(r)$. We validate the robustness of this method and the assumptions in Sec. IV of the Supplemental Material.
@@
Our analysis of the resulting potentials for S06A044 and S08A06 indeed yields one minimum and one maximum at all $t$. The locations of the respective LRs exhibit minor periodic variation in line with the BS oscillations observed in the other diagnostics, but the pair of LRs remains in place with no indication of instability. We show these variations together with selected profiles of the AEP over a representative oscillation cycle in Fig. 4 for S06A044. The same behaviour, merely with smaller oscillation amplitude, is found for S08A06.
@@
Figure 5: Instability timescales as a function of $A_{0}$ for solitonic BSs with $\sigma_{0}=0.08$. The vertical lines mark the transition between stable and unstable branches as predicted by linear theory; cf. Fig. 1. The marker style indicates dynamical fate while the color indicates the code used: black for Sph, lime green for Axi, orange for L3D. BS models whose time evolution show no sign of instability are represented with arrows at the top of the figure.
Nonlinear stability—We complement our detailed analysis of the BS models S06A044 and S08A06 with a larger set of time evolutions of the $\sigma=0.08$ BS sequence. We have used all codes for this analysis, but restricted 3D evolutions to $\mu t\lesssim 3000$ for reasons of computational cost. In this analysis, we define the instability time as the moment when $A_{0}(t)$ departs by $10\,\%$ from its equilibrium value. If $A_{0}(t)$ yields no signature of departing from the oscillatory pattern seen in Fig. 2 within $3000\,\mu^{-1}$, the model is designated stable. The result is shown in Fig. 5 and fully agrees with the perturbative predictions: models on unstable branches succumb to exponentially growing perturbations on a timescale that diverges at the threshold to stable branches. These transition points coincide with extrema in the $M(A_{0})$ curves, i.e. points where the oscillation frequency $\chi_{0}$ of the fundamental mode switches between real and imaginary.